# Circular-beam approximation for quantum channels in the turbulent atmosphere

### v1

@@ Abstract
The evolution of quantum states of light in free-space channels is strongly influenced by atmospheric turbulence, posing a significant challenge for quantum communication. The transmittance in such channels randomly fluctuates. This effect is commonly described by the probability distribution of transmittance (PDT). The elliptic-beam approximation provides an analytical model for the PDT, showing good agreement with experimental and simulation data within a specific range of channel parameters. In this work, we introduce the circular-beam approximation—a simplified alternative that offers satisfactory accuracy while significantly reducing computational complexity. We also present an alternative technique for evaluating the parameters of this model based on the first two moments of transmittance. This approach notably extends the applicability range of our PDT model, offering a practical tool for characterizing atmospheric channels in quantum applications.
@@ I Introduction
To understand how this effect influences the quantum state of light, it is first necessary to specify the method of quantum information encoding at the transmitter and its detection at the receiver. In this context, several encoding techniques can be considered, including encoding via spatial [32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43] and temporal [44] mode structures, polarization modes [5, 8, 45, 46], as well as states of a quasi-monochromatic mode in the form of a Gaussian light beam with subsequent photocounting measurements [47, 48] or homodyne detection [49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 12, 61, 62, 63, 64, 65, 66, 6, 7, 49, 12, 67], among others. The effect of atmospheric turbulence on the quantum state in the latter two encoding schemes can be characterised by random fluctuations in the transmittance efficiency: as the beam shape randomly fluctuates at the receiver, the fraction of light passing through the receiver aperture also varies accordingly. In the following, we focus our analysis on this scenario.
@@
Let the quantum states at the receiver and transmitter stations be described by the Glauber-Sudarshan $P$ functions [68, 69] $P_{\text{out}}\left(\alpha\right)$ and $P_{\text{in}}\left(\alpha\right)$, respectively. As discussed in Ref. [70] (see also a detailed derivation in Ref. [44]), these states are related to each other as
[…]
@@
Here, $\eta\in[0,1]$ is a random transmittance and $\mathcal{P}(\eta)$ is the probability distribution of transmittance (PDT)—a function characterising quantum free-space channels. Based on this input-output relation, the transmission of quantum properties through the atmosphere channel [71, 72, 45, 46, 65, 66, 73, 74] and quantum communication protocols [61, 75, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 62, 63, 64] have been analyzed.
@@
These results demonstrate that having an appropriate PDT model is crucial for the accurate description of free-space quantum channels. The simplest analytical model is given by the log-normal distribution [47, 48], which must be properly truncated to the domain $\eta\in[0,1]$. Another heuristic model, which shows a good agreement with numerical simulations, is based on the Beta distribution; see Ref. [76]. Assuming statistically independent contributions of beam wandering and beam-spot destorsion, a model based on the law of total probability was derived in Ref. [75]. The beam-wandering model, assuming a Gaussian beam profile at the receiver and considering the effect of beam wandering, was introduced in Ref. [72]. Its modification, which additionally accounts for random fluctuations of an elliptically shaped beam, was presented in Ref. [77]. The latter provides a good fit to experimental data and remains effective under diverse weather conditions [13].
@@
Nevertheless, two issues are associated with the elliptic-beam model: (i) it requires complex calculations involving approximations and numerical simulations; (ii) although it provides a good fit to experimental data over a broad range of channel parameters, numerical simulations in Ref. [76] have shown that the employed parameter evaluation method introduces a significant bias in the distribution mode, causing the model to perform reliably only within a narrow range of channel parameters. The aim of this paper is to address these issues. We propose a simplified version of the elliptic-beam approximation, assuming that the beam retains a Gaussian profile but with a circular shape of a random radius—the circular-beam approximation. While this model reduces computational complexity, it still exhibits a notable bias in the distribution mode across a significant range of parameter values. This limitation is primarily due to the strongly non-Gaussian character of the beam shape at the receiver. To overcome this, we introduce an alternative parameter evaluation method based on matching the first two moments of the transmittance distribution.
@@
The rest of the paper is organised as follows. In Sec. II, we provide the preliminary information necessary for introducing our method. The circular-beam model is presented in Sec. III. In Sec. IV, we describe a method for evaluating the PDT parameters based on matching the first two moments of the transmittance. Analytical expressions for the model parameters, derived within the framework of the Huygens–Kirchhoff phase approximation, are given in Sec. V. Examples of the PDT and validation of our models through numerical simulations are discussed in Sec. VI. In Sec. VII, we present examples demonstrating the application of our techniques to describe the transmission of nonclassical properties of quantum light through free-space channels. Finally, a summary and conclusions are provided in Sec. VIII.
@@ II Preliminaries
In this section, we provide the reader with a preliminary background on free-space channels [78, 79, 80, 81, 82], which is essential for formulating our approach. We also summarize key aspects of the beam-wandering model [72], which plays a crucial role in our technique. Our basic assumption is that quantum states of light are prepared in a quasi-monochromatic mode, representing a pulsed Gaussian beam. Since Eq. (1) holds for both quantum and classical fields, the derivation of the PDT can be carried out by analyzing the transmittance of a classical quasi-monochromatic mode through the receiver aperture; for details, see Appendix A of Ref. [76].
Let $u(\textbf{r};z)$ be the beam amplitude, where $\textbf{r}=(x,y)$ and $z$ are the transverse and axial coordinates in the propagation direction, respectively. In the paraxial approximation, this amplitude satisfies the equation (see, e.g., Ref. [80]),
$0\displaystyle 2ik\frac{\partial u(\textbf{r};z)}{\partial z}+\Delta_{\textbf{r}}u(\textbf{r};z)+2k^{2}\delta n(\textbf{r};z)u(\textbf{r};z)=$ 0 . (2)
@@
where $W_{0}$ and $F_{0}$ are the beam-spot radius and the wavefront radius at the transmitter, respectively.
@@
where $z=L$ is the axial coordinate of the aperture plane (channel length). Random fluctuations in $\delta n(\textbf{r};L)$ lead to random fluctuations in $I(\textbf{r};L)$. For the purposes of our paper, it is sufficient to consider only three random variables that characterize this random field. The first variable is the random transmittance efficiency defined as
@@
$\Gamma_{2}(\textbf{r};z)=\left\langle I(\textbf{r};z)\right\rangle$ (10)
@@
$\Gamma_{4}(\textbf{r}_{1},\textbf{r}_{2};z)=\left\langle I(\textbf{r}_{1};z)I(\textbf{r}_{2};z)\right\rangle,$ (11)
respectively. These functions can be obtained using methods of classical atmospheric optics; see, e.g., discussions in Sec. V. The first and second moments of the transmittance are given by
@@
$=\left\langle x_{0}^{2}\right\rangle=\int_{\mathds{R}^{4}}d^{2}\textbf{r}_{1}d^{2}\textbf{r}_{2}x_{1}x_{2}\Gamma_{4}(\textbf{r}_{1},\textbf{r}_{2};z).$
The mean squared beam-spot radius reads
$\left\langle S\right\rangle=4\left(\int_{\mathds{R}^{2}}d^{2}\textbf{r}x^{2}\Gamma_{2}(\textbf{r};z)-\left\langle x_{0}^{2}\right\rangle\right).$ (15)
The second moment of this random variable requires a more involved analysis, which is presented in Sec. III.
In the circular-beam approximation presented in this paper, we assume that the instant intensity has a Gaussian form,
$I(\textbf{r},z)=\frac{2}{\pi S}\exp\left(-\frac{2}{S}|\textbf{r}-\textbf{r}_{0}|^{2}\right).$ (16)
Following a procedure similar to that used for the beam-wandering model [72], and assuming that the transmittance fluctuations arise solely from centroid fluctuations, we obtain the conditional PDT for a fixed value of $S$,
@@
×$\times\exp\left[-\frac{R^{2}(S)}{2\sigma^{2}_{\text{bw}}}\left(\ln\frac{\eta_{0}(S)}{\eta}\right)^{2/\lambda(S)}\right]$
@@
×$\times\left[\ln\left(\frac{2\eta_{0}(S)}{1-\exp\left(-4\frac{a^{2}}{S}\right)\operatorname{I}_{0}\left(4\frac{a^{2}}{S}\right)}\right)\right]^{-1}$ (20)
@@
$R(S)=a\left[\ln\left(\frac{2\eta_{0}(S)}{1-\exp\left(-4\frac{a^{2}}{S}\right)\operatorname{I}_{0}\left(4\frac{a^{2}}{S}\right)}\right)\right]^{-1/\lambda(S)}.$ (21)
is the scale parameter, $a$ is the aperture radius, $\operatorname{I}_{n}(x)$ is the modified Bessel function.
By setting $S=\left\langle S\right\rangle$ in Eq. (17), we recover the beam-wandering PDT [72], which neglects fluctuations in the beam-spot radius. A more realistic model, the elliptic-beam approximation [77], still assumes a Gaussian beam profile but with an elliptical shape. This approach involves a more complex analysis due to the increased number of fluctuating parameters. In this work, we aim to simplify the description by assuming that the beam retains a circular Gaussian shape, while allowing the squared beam-spot radius $S$ to fluctuate.
@@ III Circular-beam approximation
In this section, we introduce the main concept of the circular-beam approximation and describe a method for determining its parameters by matching the first two moments of the squared beam-spot radius $S$. This approach serves as a straightforward simplification of the elliptic-beam approximation presented in Ref. [77]. As a result, it avoids the need for involved analytical and numerical calculations used in the elliptic-beam approximation.
@@
[…]
As in the elliptic-beam model, we assume for the circular-beam model that $S$ follows a log-normal distribution [83],
@@
Our first task is to validate the assumption that the probability distribution for $S$ can be good approximated by the log-normal distribution (23). To this end, we simulate the beam propagation through turbulent atmosphere by using the sparse-spectrum model [84, 85, 86] for the phase-screen method [87, 88, 89, 90]. Following the procedure detailed in Ref. [76], we obtain a sample $\{S_{i}|i=1\ldots N\}$ of $N$ realizations of the squared beam-spot radius $S$. This sample is then used to estimate the parameters $\mu$ and $\sigma^{2}$ of the log-normal distribution. The resulting parameters define the log-normal distribution, which we compare with the sampled distribution of $S$. In particular, we directly reconstruct $P(S)$ by the smooth-kernel method and compare it with the related log-normal distribution. Additionally, we estimate the empirical cumulative probability distribution function,
@@
where $\theta(S-S_{i})$ is the Heaviside step-function, and compare it to the cumulative distribution function of the log-normal distribution,
[…]
using the Kolmogorov-Smirnov statistics,
@@
Here $\operatorname{erfc}(x)$ denotes the complementary error function. Our results show good agreement between the sampled data and the log-normal distribution for channels with weak impact of turbulence, i.e. with the Rytov parameter
$\sigma_{R}^{2}=1.23C_{n}^{2}k^{7/6}L^{11/6}$ (27)
@@
Figure 1: Comparison between the log-normal and simulated probability distributions of $S$. (a): Solid and dashed lines show the probability distribution functions $P(S)$ for the log-normal distribution (23) and the distribution estimated from the simulated data, respectively. (b): The Kolmogorov-Smirnov statistics (26) as a function of the Rytov parameter. Here $L$ ranges from $500$ m to $5000$ m, $\lambda=808$ nm, $W_{0}=\sqrt{L\lambda/\pi}$, $C_{n}^{2}=10^{-15}$ m-2/3, $l_{0}=10^{-6}$ m, $L_{0}=5\times 10^{3}$ m, $F_{0}=L$.
Our next task consists in formulation of methods for evaluating three parameters of the circular-beam approximation based on the field correlation functions $\Gamma_{2}(\textbf{r};z)$ and $\Gamma_{4}(\textbf{r}_{1},\textbf{r}_{2};z)$. The variance of the beam-centroid coordinate, $\sigma^{2}_{\text{bw}}$, is given by Eq. (14). The parameters of the log-normal distribution are expressed in terms of the first two moments, $\left\langle S\right\rangle$ and $\left\langle S^{2}\right\rangle$, as
For deriving $\left\langle S^{2}\right\rangle$, we start by squaring and averaging the instant beam-spot radius (9),
$\left\langle S^{2}\right\rangle=16\bigg{(}\int_{\mathds{R}^{4}}d^{2}\textbf{r}_{1}d^{2}\textbf{r}_{2}x_{1}^{2}x_{2}^{2}\Gamma_{4}(\textbf{r}_{1},\textbf{r}_{2};z)$ (30)
$-2\left\langle x_{0}^{2}\int_{\mathds{R}^{2}}d^{2}\textbf{r}x^{2}I(\textbf{r};z)\right\rangle+\left\langle x_{0}^{4}\right\rangle\bigg{)}.$
$\left\langle x_{0}^{2}\int_{\mathds{R}^{2}}d^{2}\textbf{r}x^{2}I(\textbf{r};z)\right\rangle=\frac{1}{4}\left\langle x_{0}^{2}S\right\rangle+\left\langle x_{0}^{4}\right\rangle.$ (31)
see Fig. 2. The absolute value of the coefficient remains significantly below unity, thereby justifying the assumption of weak correlation. This yields,
Combining all these assumptions in Eq. (30), we arrive at its approximate form
$\left\langle S^{2}\right\rangle\approx 16\bigg{(}\int_{\mathds{R}^{4}}d^{2}\textbf{r}_{1}d^{2}\textbf{r}_{2}x_{1}^{2}x_{2}^{2}\Gamma_{4}(\textbf{r}_{1},\textbf{r}_{2};z)$ (35)
Figure 2: The Pearson correlation coefficient between $x_{0}^{2}$ and $S$, cf. Eq. (32), as a function of the Rytov parameter. For the rest channel parameters, see Fig. 1.
Therefore, the application of the circular-beam approximation can be outlined as follows. First, we calculate the parameters $\sigma_{\mathrm{bw}}^{2}$, $\left\langle S\right\rangle$, and $\left\langle S^{2}\right\rangle$ using Eqs. (14), (15), and (35), respectively. Second, we substitute the obtained values of $\left\langle S\right\rangle$ and $\left\langle S^{2}\right\rangle$ into Eqs. (28) and (29) to determine the log-normal parameters $\mu$ and $\sigma$. Finally, these parameters are used in Eqs. (22) and (23) to calculate the PDT.
IV Matching transmittance moments
@@
As discussed in Ref. [76], the beam-wandering and elliptic-beam approximations satisfactorily model the shapes of the PDT when compared with numerical simulations. However, the modes of these distributions are significantly biased over a wide range of aperture radii $a$. Another problem is that beyond this domain of validity, the first two moments of the transmittance, $\left\langle\eta\right\rangle$ and $\left\langle\eta^{2}\right\rangle$, are also biased. Naturally, this issue is also inherent in the circular-beam approximation presented in the previous section. In this section, we address this problem by introducing an alternative method for evaluating the log-normal parameters $\mu$ and $\sigma^{2}$ in Eq. (23). Specifically, instead of matching them to the first two moments of the squared beam-spot radius, we match them to the first two moments of the transmittance, $\left\langle\eta\right\rangle$ and $\left\langle\eta^{2}\right\rangle$, given by Eqs. (12) and (13), respectively.
The method is based on the fact that the first two moments of the circular-beam PDT (22) are given by
[…]
@@
$\eta\right>_{S}=1-\exp\!\left(-2\frac{a^{2}}{4\left<x_{0}^{2}\right>+S}\right),$ (38)
$\eta^{2}\right>_{S}$ $=1-2\exp\!\left(-2\frac{a^{2}}{4\left<x_{0}^{2}\right>+S}\right)$ (39)
@@
$\beta=(2p+1)^{-1},\quad p=\frac{1}{8}\frac{S}{\left<x_{0}^{2}\right>},$ (41)
and $Q(x,y)$ is the Marcum $Q$-function of the first order; see Refs. [92, 93, 94].
@@
The transmittance moment matching method for the circular-beam approximation can now be summarized as follows. First, we calculate the parameters $\sigma_{\mathrm{bw}}^{2}$, $\langle\eta\rangle$, and $\langle\eta^{2}\rangle$ using Eqs. (14), (12), and (13), respectively. Second, we treat Eqs. (36) and (37) as a system of algebraic equations for $\mu$ and $\sigma^{2}$, and solve it numerically via least-squares optimization. As an initial guess one can use the values of $\mu$ and $\sigma^{2}$ obtained with the method of matching the moments of $S$, described in the previous section [^1] In practice, we also constrain the values of $\mu$ and $\sigma^{2}$. Specifically, we require the mean of the corresponding $S$ distribution to lie within a factor of five of its initial guess. Additionally, we require that the probability of an instantaneous value of $S$ exceeding 10 times its mean is less than 1%. For the considered case, this leads to the constraint $210^{-6}<\sigma^{2}<$ 2. The resulting values of $\mu$ and $\sigma^{2}$ are then used in Eqs. (22) and (23) to evaluate the PDT. Since $\langle\eta\rangle$ and $\langle\eta^{2}\rangle$ are now directly matched, their values are no longer biased.
@@
In realistic applications, constant (non-fluctuating) losses—caused by absorption, scattering, and imperfections in the transmitter and receiver optical systems—are always present in addition to the fluctuating losses induced by atmospheric turbulence. These constant losses may be naturally incorporated into the transmittance moment matching method. To do so, one replaces the left-hand sides of Eqs. (36) and (37) with $\eta_{\mathrm{c}}\langle\eta\rangle$ and $\eta_{\mathrm{c}}^{2}\langle\eta^{2}\rangle$, respectively, where $\eta_{\mathrm{c}}$ represents the efficiency associated with the constant losses. In other words, the modified left-hand sides of these equations correspond to realistic transmittance moments that account for all experimental losses. Within this framework, the constant losses are effectively modeled by adjusting the parameters $\mu$ and $\sigma^{2}$ of the log-normal distribution for $S$. This approach represents an optional modification of the described method, applicable in suitable scenarios.
V Phase approximation of the Huygens-Kirchhoff method

### v2

# Circular-beam approximation for quantum channels in a turbulent atmosphere

@@ Abstract
The evolution of quantum states of light in free-space channels is strongly influenced by atmospheric turbulence, posing a significant challenge for quantum communication. The transmittance in such channels randomly fluctuates. This effect is commonly described by the probability distribution of transmittance (PDT). The elliptic-beam approximation provides an analytical model for the PDT, showing good agreement with experimental and simulation data within a specific range of channel parameters. In this work, we introduce the circular-beam approximation—a simplified alternative that offers satisfactory accuracy while significantly reducing computational complexity. Our method naturally leads to a technique for determining the model parameters from the first two moments of the transmittance. This approach eliminates the model misspecification bias inherent in the elliptic-beam approximation and significantly extends the applicability range of the PDT model, providing a practical tool for characterizing atmospheric channels in quantum applications.
@@ I Introduction
To understand how this effect influences the quantum state of light, it is first necessary to specify the method of quantum information encoding at the transmitter and its detection at the receiver. In this context, several encoding techniques can be considered, including encoding via spatial [32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43] and temporal [44] mode structures, polarization modes [5, 8, 45, 46], as well as states of a quasimonochromatic mode in the form of a Gaussian light beam with subsequent photocounting measurements [47, 48] or homodyne detection [49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 12, 61, 62, 63, 64, 65, 66, 6, 7, 49, 12, 67], among others. The effect of atmospheric turbulence on the quantum state in the latter two encoding schemes can be characterized by random fluctuations in the transmittance efficiency: as the beam shape randomly fluctuates at the receiver, the fraction of light passing through the receiver aperture also varies accordingly. In the following, we focus our analysis on this scenario.
@@
Let the quantum states at the receiver and transmitter stations be described by the Glauber–Sudarshan $P$ functions [68, 69] $P_{\text{out}}\left(\alpha\right)$ and $P_{\text{in}}\left(\alpha\right)$, respectively. As discussed in Ref. [70] (see also a detailed derivation in Ref. [44]), these states are related to each other as
$P_{\text{out}}\left(\alpha\right)=\int_{0}^{1}d\eta\mathcal{P}(\eta)\frac{1}{\eta}P_{\text{in}}\left(\frac{\alpha}{\sqrt{\eta}}\right).$ (1)
@@
Here, $\eta\in[0,1]$ is a random transmittance and $\mathcal{P}(\eta)$ is the probability distribution of transmittance (PDT)—a function characterizing quantum free-space channels. Based on this input–output relation, the transmission of quantum properties through atmosphere channels [71, 72, 45, 46, 65, 66, 73, 74] and quantum communication protocols [61, 75, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 62, 63, 64] has been analyzed.
@@
These results demonstrate that having an appropriate PDT model is crucial for the accurate description of free-space quantum channels. The simplest analytical model is given by the log-normal distribution [47, 48], which must be properly truncated to the domain $\eta\in[0,1]$. Another heuristic model, which shows a good agreement with numerical simulations, is based on the beta distribution; see Ref. [76]. Assuming statistically independent contributions of beam–wandering and beam-spot distorsion, a model based on the law of total probability was derived in Ref. [75]. The beam-wandering model, assuming a Gaussian beam profile at the receiver and considering the effect of beam wandering, was introduced in Ref. [72]. Its modification, which additionally accounts for random fluctuations of an elliptically shaped beam, was presented in Ref. [77]. The latter provides a good fit to experimental data and remains effective under diverse weather conditions [13].
@@
Nevertheless, two issues are associated with the elliptic-beam model: (i) it requires complex calculations involving approximations and numerical simulations; (ii) although it provides a good fit to experimental data over a broad range of channel parameters, numerical simulations in Ref. [76] have shown that the employed parameter evaluation method introduces a significant bias in the distribution mode, causing the model to perform reliably only within a narrow range of channel parameters. The aim of this paper is to address these issues. We propose a simplified version of the elliptic-beam approximation, assuming that the beam retains a Gaussian profile but with a circular shape of a random radius—the circular-beam approximation. While this model reduces computational complexity, it still exhibits a notable bias in the distribution mode across a significant range of parameter values. This limitation is primarily due to the strongly non-Gaussian character of the beam shape at the receiver. To overcome this model–misspecification bias, we introduce an alternative parameter evaluation method based on matching the first two moments of the transmittance distribution.
@@
The rest of the paper is organized as follows. In Sec. II we provide the preliminary information necessary for introducing our method. Section III presents the circular-beam model and the associated method for determining the PDT parameters, which eliminates model misspecification bias. Analytical expressions for the model parameters, derived within the framework of the Huygens–Kirchhoff phase approximation, are given in Sec. IV. Examples of the PDT and validation of our models through numerical simulations are discussed in Sec. V. In Sec. VI we present examples demonstrating the application of our techniques to describe the transmission of nonclassical properties of quantum light through free-space channels. Finally, a summary and conclusions are provided in Sec. VII.
@@ II Preliminaries
In this section we provide the reader with a preliminary background on free-space channels [78, 79, 80, 81, 82], which is essential for formulating our approach. We also summarize key aspects of the beam-wandering model [72], which plays a crucial role in our technique. Our basic assumption is that quantum states of light are prepared in a quasi-monochromatic mode, representing a pulsed Gaussian beam. Since Eq. (1) holds for both quantum and classical fields, the derivation of the PDT can be carried out by analyzing the transmittance of a classical quasi-monochromatic mode through the receiver aperture; for details, see Appendix A of Ref. [76].
Let $u(\textbf{r};z)$ be the beam amplitude, where $\textbf{r}=(x,y)$ and $z$ are the transverse and axial coordinates in the propagation direction, respectively. In the paraxial approximation, this amplitude satisfies the equation (see, e.g., Ref. [80])
$2ik\frac{\partial u(\textbf{r};z)}{\partial z}+\Delta_{\textbf{r}}u(\textbf{r};z)+2k^{2}\delta n(\textbf{r};z)u(\textbf{r};z)=0.$ (2)
@@
where $W_{0}$ and $F_{0}$ are the beam-spot radius and the wave–front radius at the transmitter, respectively.
@@
where $z=L$ is the axial coordinate of the aperture plane (channel length). Random fluctuations in $\delta n(\textbf{r};L)$ lead to random fluctuations in $I(\textbf{r};L)$. For the purposes of our paper, it is sufficient to consider only three random variables that characterize this random field. The first variable is the random transmittance efficiency, defined as
@@
$\Gamma_{2}(\textbf{r};L)=\left\langle I(\textbf{r};L)\right\rangle$ (10)
@@
$\Gamma_{4}(\textbf{r}_{1},\textbf{r}_{2};L)=\left\langle I(\textbf{r}_{1};L)I(\textbf{r}_{2};L)\right\rangle,$ (11)
respectively. These functions can be obtained using methods of classical atmospheric optics; see, e.g., discussions in Sec. IV. The first and second moments of the transmittance are given by
@@
$=\left\langle x_{0}^{2}\right\rangle=\int_{\mathds{R}^{4}}d^{2}\textbf{r}_{1}d^{2}\textbf{r}_{2}x_{1}x_{2}\Gamma_{4}(\textbf{r}_{1},\textbf{r}_{2};L).$
The mean–squared beam-spot radius reads
$\left\langle S\right\rangle=4\left(\int_{\mathds{R}^{2}}d^{2}\textbf{r}x^{2}\Gamma_{2}(\textbf{r};L)-\left\langle x_{0}^{2}\right\rangle\right).$ (15)
The second moment of this random variable requires a more involved analysis, which is presented in Appendix A.
In the circular-beam approximation presented in this paper, we assume that the instant intensity has a Gaussian form:
$I(\textbf{r},L)=\frac{2}{\pi S}\exp\left(-\frac{2}{S}|\textbf{r}-\textbf{r}_{0}|^{2}\right).$ (16)
Following a procedure similar to that used for the beam-wandering model [72] and assuming that the transmittance fluctuations arise solely from centroid fluctuations, we obtain the conditional PDT for a fixed value of $S$,
@@
$\times\exp\left[-\frac{R^{2}(S)}{2\sigma^{2}_{\text{bw}}}\left(\ln\frac{\eta_{0}(S)}{\eta}\right)^{2/\lambda(S)}\right]$
@@
$\times\left[\ln\left(\frac{2\eta_{0}(S)}{1-\exp\left(-4\frac{a^{2}}{S}\right)\operatorname{I}_{0}\left(4\frac{a^{2}}{S}\right)}\right)\right]^{-1}$ (20)
@@
$R(S)=a\left[\ln\left(\frac{2\eta_{0}(S)}{1-\exp\left(-4\frac{a^{2}}{S}\right)\operatorname{I}_{0}\left(4\frac{a^{2}}{S}\right)}\right)\right]^{-1/\lambda(S)}$ (21)
is the scale parameter, $a$ is the aperture radius, and $\operatorname{I}_{n}(x)$ is the modified Bessel function.
By setting $S=\left\langle S\right\rangle$ in Eq. (17), we recover the beam-wandering PDT [72], which neglects fluctuations in the beam-spot radius. A more realistic model, the elliptic-beam approximation [77], still assumes a Gaussian beam profile but with an elliptical shape. This approach involves a more complex analysis due to the increased number of fluctuating parameters. In this work we aim to simplify the description by assuming that the beam retains a circular Gaussian shape while allowing the squared beam-spot radius $S$ to fluctuate.
@@ III Circular-beam approximation
In this section we introduce the main concept of the circular-beam approximation and describe a method for determining its parameters by matching the first two moments of the transmittance $\eta$. On the one hand, this approach provides a straightforward simplification of the elliptic-beam approximation presented in Ref. [77], thereby avoiding the need for the more involved analytical and numerical calculations required in that model. On another hand, this simplification enables a direct matching of the model parameters to the transmittance moments, thus reducing the model misspecification bias that arises in the elliptic-beam approximation over a wide range of aperture radii $a$, see Ref. [76].
@@
$\mathcal{P}(\eta)=\int_{0}^{\infty}dS\mathcal{P}(\eta|S)P(S).$ (22)
As in the elliptic-beam model, we assume for the circular-beam model that $S$ follows a log-normal distribution [83]:
@@
Our first task is to validate the assumption that the probability distribution for $S$ can be well approximated by the log-normal distribution (23). To this end, we simulate the beam propagation through turbulent atmosphere by using the sparse-spectrum model [84, 85, 86] for the phase-screen method [87, 88, 89, 90]. Following the procedure detailed in Ref. [76], we obtain a sample $\{S_{i}|i=1\ldots N\}$ of $N$ realizations of the squared beam-spot radius $S$. This sample is then used to estimate the parameters $\mu$ and $\sigma^{2}$ of the log-normal distribution. The resulting parameters define the log-normal distribution, which we compare with the sampled distribution of $S$. In particular, we directly reconstruct $P(S)$ by the smooth-kernel method and compare it with the related log-normal distribution. Additionally, we estimate the empirical cumulative probability distribution function,
@@
where $\theta(S-S_{i})$ is the Heaviside step function, and compare it to the cumulative distribution function of the log-normal distribution,
$F(S)=\int_{0}^{S}dS^{\prime}P(S^{\prime})=\frac{1}{2}\operatorname{erfc}\left(\frac{\ln S-\mu}{\sigma\sqrt{2}}\right),$ (25)
using the Kolmogorov–Smirnov statistics,
@@
Here $\operatorname{erfc}(x)$ denotes the complementary error function. Our results show good agreement between the sampled data and the log-normal distribution for channels with weak impact of turbulence, i.e. with the Rytov parameter,
$\sigma_{R}^{2}=1.23C_{n}^{2}k^{7/6}L^{11/6},$ (27)
@@
Figure 1: Comparison between the log-normal and simulated probability distributions of $S$. (a) Solid and dashed lines show the probability distribution functions $P(S)$ for the log-normal distribution (23) and the distribution estimated from the simulated data, respectively. (b) The Kolmogorov–Smirnov statistics (26) as a function of the Rytov parameter. Here $L$ ranges from $500$ to $5000$ m, $\lambda=808$ nm, $W_{0}=\sqrt{L\lambda/\pi}$, $C_{n}^{2}=10^{-15}$ m-2/3, $l_{0}=10^{-6}$ m, $L_{0}=5\times 10^{3}$ m, $F_{0}=L$.
Our next task consists in formulation of methods for evaluating three parameters of the circular-beam approximation based on the field correlation functions $\Gamma_{2}(\textbf{r};L)$ and $\Gamma_{4}(\textbf{r}_{1},\textbf{r}_{2};L)$. While the beam-centroid coordinate $\sigma^{2}_{\text{bw}}$ can be determined from Eq. (14), the methods for determining the other two parameters, $\mu$ and $\sigma^{2}$, must be formulated separately. A straightforward approach is to match them to the moments of the squared beam-spot radius, $\langle S\rangle$ and $\langle S^{2}\rangle$, as outlined in Appendix A. However, as shown in Ref. [76] for the elliptic-beam approximation, this approach leads to a pronounced shift of the PDT mode (its maximum) across a wide range of aperture radii $a$. Another problem is that beyond a limited domain of validity, the first two moments of the transmittance, $\left\langle\eta\right\rangle$ and $\left\langle\eta^{2}\right\rangle$, are also biased. The reason for this is caused by the fact that in realistic scenarios the beam profile is strongly non-Gaussian. This implies that the beam-wandering, circular-beam, and elliptic-beam approximations are all affected by model misspecification bias.
@@
In this section we address this problem by introducing an alternative method for evaluating the log-normal parameters $\mu$ and $\sigma^{2}$ in Eq. (23). Specifically, instead of matching them to the first two moments of the squared beam-spot radius, we match them to the first two moments of the transmittance, $\left\langle\eta\right\rangle$ and $\left\langle\eta^{2}\right\rangle$, given by Eqs. (12) and (13), respectively. The method is based on the fact that the first two moments of the circular-beam PDT (22) are given by
$\left\langle\eta\right\rangle=\int_{0}^{+\infty}dSP(S|\mu,\sigma^{2})\left\langle\eta\right\rangle_{S},$ (28)
$\left\langle\eta^{2}\right\rangle=\int_{0}^{+\infty}dSP(S|\mu,\sigma^{2})\left\langle\eta^{2}\right\rangle_{S}.$ (29)
@@
$\left<\eta\right>_{S}=1-\exp\!\left(-2\frac{a^{2}}{4\sigma^{2}_{\text{bw}}+S}\right),$ (30)
$\left<\eta^{2}\right>_{S}$ $=1-2\exp\!\left(-2\frac{a^{2}}{4\sigma^{2}_{\text{bw}}+S}\right)$ (31)
@@
$\beta=(2p+1)^{-1},\quad p=\frac{1}{8}\frac{S}{\sigma^{2}_{\text{bw}}},$ (33)
and $Q(x,y)$ is the Marcum $Q$ function of the first order; see Refs. [92, 93, 94].
@@
The transmittance–moment matching method for the circular-beam approximation can now be summarized as follows. First, we calculate the parameters $\sigma_{\mathrm{bw}}^{2}$, $\langle\eta\rangle$, and $\langle\eta^{2}\rangle$ using Eqs. (14), (12), and (13), respectively. Second, we treat Eqs. (28) and (29) as a system of algebraic equations for $\mu$ and $\sigma^{2}$, and solve it numerically via least-squares optimization. As an initial guess one can use the values of $\mu$ and $\sigma^{2}$ used in the method of matching the moments of $S$, described in Appendix A; see Eqs. (49) and (50). [^1] In practice, we also constrain the values of $\mu$ and $\sigma^{2}$. Specifically, we require the mean of the corresponding $S$ distribution to lie within a factor of 5 of its initial guess. Additionally, we require that the probability of an instantaneous value of $S$ exceeding 10 times its mean is less than 1%. For the considered case, this leads to the constraint $10^{-6}<\sigma^{2}<2$. The resulting values of $\mu$ and $\sigma^{2}$ are then used in Eqs. (22) and (23) to evaluate the PDT. Since $\langle\eta\rangle$ and $\langle\eta^{2}\rangle$ are now directly matched, their values are no longer biased.
@@
In realistic applications, constant (nonfluctuating) losses—caused by absorption, scattering, and imperfections in the transmitter and receiver optical systems—are always present, in addition to the fluctuating losses induced by atmospheric turbulence. These constant losses may be naturally incorporated into the transmittance–moment matching method. To do so, one replaces the left-hand sides of Eqs. (28) and (29) with $\eta_{\mathrm{c}}\langle\eta\rangle$ and $\eta_{\mathrm{c}}^{2}\langle\eta^{2}\rangle$, respectively, where $\eta_{\mathrm{c}}$ represents the efficiency associated with the constant losses. In other words, the modified left-hand sides of these equations correspond to realistic transmittance moments that account for all experimental losses. Within this framework, the constant losses are effectively modeled by adjusting the parameters $\mu$ and $\sigma^{2}$ of the log-normal distribution for $S$. This approach represents an optional modification of the described method, applicable in suitable scenarios.
IV Phase approximation of the Huygens–Kirchhoff method